OpenStudy (anonymous):

prove that R* x R is a group under the operation * defined by (a,b) * (c,d) = (ac, bc + d)

OpenStudy (anonymous):

is it \((a, b)\circ (c,d)=(ac, bc+d)\) ?

OpenStudy (anonymous):

and i guess we need to know what \(R^*\) means in this context i take it \(R\) is a ring?

OpenStudy (anonymous):

R is all real numbers

OpenStudy (anonymous):

R* is a multiplicative group of all real numbers

OpenStudy (amistre64):

R* is the nonzero reals

OpenStudy (anonymous):

i think the biggest challenge will be to find the inverse first figure out what the identity is

OpenStudy (amistre64):

(a,b) * (c,d) = (ac, bc + d) (ac, bc + d) * (x,y) = (acx, bcx+dx+y) is this associative?

OpenStudy (anonymous):

i hate checking associativity, it is a pain used to get away with saying "in inherits it from the original group" but not in this case

OpenStudy (amistre64):

(c,d) * (x,y) = (cx, dx + y) (a,b) * (cx, dx + y) = (acx, bcx + dx+y) yay! its associable

OpenStudy (anonymous):

whew

OpenStudy (amistre64):

(a,b) * (x,y) = (a,b) is identity

OpenStudy (anonymous):

now for the identity, solve \[(x,y)(c,d)=(c,d)\] for \((x,y)\) you get \[(x,y)(c,d)=(xc,yc+d)=(c,d)\implies xc=c\implies x=1\] and \[yc+d=d\implies y=0\]

OpenStudy (anonymous):

identity is \((1,0)\) now for inverses

OpenStudy (amistre64):

can we assume commutability in groups?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but we could check to see if it is commutative i guess hmmm

OpenStudy (anonymous):

oh seems unlikely since it is not symmetric in the second coordinate

OpenStudy (amistre64):

and identity elements needs to be unique, correct?

OpenStudy (anonymous):

biggest job i think it to find \((a,b)^{-1}\)

OpenStudy (anonymous):

yeah identity is always unique

OpenStudy (anonymous):

doesnt it only have to be commutative for abelian groups?

OpenStudy (amistre64):

commutative is the definition of abelian yes

OpenStudy (anonymous):

i doubt this is commutative you still have to show it contains inverses

OpenStudy (anonymous):

I mean from what you have already done, haven't you proved that is it a group?

OpenStudy (anonymous):

no amistre showed it was associative i found the identity element you still have to show it contains inverses

OpenStudy (anonymous):

and i guess you have to show it is closed under the operation too

OpenStudy (amistre64):

(a,b)*(1,0) = (a1,b1+0) = (a,b) (1,0)*(a,b) = (1a, 0a + b) = (a,b)

OpenStudy (amistre64):

since R*xR does not contain the (0,n) element, we would have to see of this is possible

OpenStudy (amistre64):

well, since 0 is not a first element in any of the cases, its closes

OpenStudy (anonymous):

guess this is a good time to review the axioms 1) set must be closed under the operation 2) the operation must be associative 3) there must exist and identity element \(e\) such that \(ex=xe=x\) for all \(x\in G\) 4) it must contain inverses, i.e. if \(x\in G\) there must exist an \(x^{-1}\) with \[x^{-1}x=xx^{-1}=e\]

OpenStudy (anonymous):

we got 1 through 3, now need 4

OpenStudy (amistre64):

(a,b)*(x,y) = (1,0) :)

OpenStudy (amistre64):

since we got reals and no zero firsts, ax = 1 is possible

OpenStudy (amistre64):

bx+y = 0 when x=-y/b is possible

OpenStudy (amistre64):

hmm, for b not= 0 tho

OpenStudy (amistre64):

theres no reason why y=-bx is not possible tho

OpenStudy (anonymous):

in fact it is \(y=-\frac{b}{a}\) i think

OpenStudy (amistre64):

\[(a,b)*(\frac1a,-\frac ba)=(\frac aa,\frac ba-\frac ba)=(1,0)\]

OpenStudy (amistre64):

sin ce a in R*, a not= 0

OpenStudy (amistre64):

is that commutative? :)

OpenStudy (anonymous):

\[(a,b)(x,y)=(ax,bx+y)=(1,0)\] \[ax=1\implies x=\frac{1}{a}\] \[bx+y=0\implies \frac{b}{a}+y=0\implies y=-\frac{b}{a}\] so inverse of \((a,b)\) is \[(a,b)^{-1}=(\frac{1}{a},-\frac{b}{a})\]

OpenStudy (anonymous):

oh i am a little behind i see

OpenStudy (amistre64):

must be the weather up there

OpenStudy (amistre64):

the inverse as also communicable

OpenStudy (anonymous):

still g od da mn cold!!

OpenStudy (amistre64):

been nice here in florida :)

OpenStudy (anonymous):

we had snow yesterday...

OpenStudy (anonymous):

yeah will i will take to you in july when it is 102

OpenStudy (amistre64):

lol, 102 is just a calm pleasant day :)

OpenStudy (anonymous):

i appreciate all the help today! get excited for next week when I study for finals. ;)

OpenStudy (anonymous):

well this was a fun exercise, time to do some work later all

OpenStudy (amistre64):

my teacher gave me a real analysis and an abstract algebra book to read over the summer

OpenStudy (anonymous):

I'm taking analysis 1 and 2 next year. get excited for that too

OpenStudy (anonymous):

@urbanderivative i hope it is clear what you have to do to solve this write down the axioms check that it satisfies each one

OpenStudy (anonymous):

@amistre64 what algebra book?

OpenStudy (amistre64):

.... something old, prolly from when she was in college 20 years ago. i cant recall the title at the moment

OpenStudy (anonymous):

so twenty years ago is old??

OpenStudy (amistre64):

:) at todays ages, yes

OpenStudy (anonymous):

if you can find a copy of "herstein" topics in algebra, i think that is still the gold standard, although i guess it is really really old

OpenStudy (anonymous):

maybe from the 60's although how we call that "mid century"

OpenStudy (anonymous):

or 'mad men era'